You have $5,436.60 in an account that pays 9.40% interest, compounded continuously. If you deposited some funds 12 years ago, how much was your original deposit? Round your answer to the nearest dollar.

a. $1,637

b. $1,566

c. $1,461

d. $1,760

e. $2,006
Answer #1

We use the formula:
A=P(e)^rn
where
A=future value
P=present value
r=rate of interest
n=time period.

e=2.71828

5,436.60=P*(2.71828)^(0.094*12)

P=5,436.60/(2.71828)^(0.094*12)

=5,436.60*0.323680215

=$1760(Approx).